# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v2

@@ 3. Deriving Epsilon based on the PRI
In this section, we describe an algorithm to choose $\epsilon$ given controllers’ privacy preferences in terms of PRIs, thus addressing Challenge 1 (Sections 3.1, 3.2). In Section 3.3, we discuss how to bound the total privacy loss in the presence of multiple queries. In Section 3.4, we provide the end-to-end DP guarantee for the algorithm. We discuss the algorithm’s complexity and extensions in Section 3.5.
@@ 3.1. Using PRI to Express Privacy Preferences
Traditionally, controllers need to choose $\epsilon$ to convey their privacy preferences. However, choosing $\epsilon$ is challenging because it is hard to understand the specific effect of $\epsilon$ on the privacy of any one individual in the dataset. In contrast, the per-individual PRI indicates the relative privacy loss of an individual compared to other individuals in the dataset. The consequence is the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output. Next, we convey the intuition on how the PRI is useful to controllers through a vignette.
Figure 1. Example of using PRI to find $\epsilon$. For illustrative purposes, we show the PRI of each student (with the corresponding query output) and leave out students’ actual GPA.
Vignette illustrating one use of the PRI. Figure 1 shows the resulting PRIs of four students after answering a query about their average GPA. If the query output were released without using DP (a), the analyst observing the output would learn a lot about those students’ grades, as indicated by the PRI (remember that the PRI is a proxy to privacy loss). If the controller used DP, they would have to choose $\epsilon$ and obtain the query output by applying a DP mechanism. This would lead to a noisy output and a PRI associated with each student (note that running the mechanism two times with the same $\epsilon$ does not guarantee the same PRIs are obtained). We observe that decreasing $\epsilon$ (b, c, d) results in noisier output, which causes the PRIs to become more similar across students. In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others. Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.2) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve. Lastly, the PRIs in both (c) and (d) indicate no student will be at a higher risk than others, but (d) produces a query output with worse accuracy. The controller may choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
The above is an illustrative example: finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
@@ 3.2. The Find-$\epsilon$-from-PRI Algorithm
When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the relative privacy loss of each individual compared to others. Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a set of PRIs that represents their preferences. We introduce an algorithm that implements this intuition, named Find-$\epsilon$-from-PRI. In Section 3.1, we have shown one criterion for choosing $\epsilon$ that leads to exactly equal PRIs, which is typically infeasible. Instead, we propose an alternative criterion to let controllers decide whether a query output respects their privacy preferences. We explain the key idea and then the algorithm.
Key Idea. We let controllers express their privacy preferences by indicating a cutoff for the highest privacy loss of any individual in the dataset. This is done by requesting exactly one parameter from the controllers, a percentage p ($0\leq p\leq 100$). We compute the minimum and maximum PRI, $PRI_{min}$ and $PRI_{max}$. Then we find an $\epsilon$ that produces a list of PRIs (and associated DP output) such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$. In other words, we let controllers specify that the maximum PRI should deviate no more than $p$ percent from the minimum PRI. For instance, setting $p=0$ means the PRIs should be exactly equal.
Unlike $\epsilon$, the parameter $p$ directly conveys the level of risk conferred to individuals after releasing a specific query output. Smaller $p$ means more privacy at the cost of worse accuracy, hence controllers can trade privacy for accuracy by setting a larger $p$. The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$.
Input : Dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$
Output : Chosen privacy parameter $\epsilon$, query output $r$
1 /* Initialize set of $\epsilon$ values to test */
@@ 2 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
3 /* Compute $q(x_{-i})$ for all individual $i$ using memoization */
@@ 7             $C[x_{i}]\leftarrow q(x_{-i})$
11       $r\leftarrow\mathcal{M}_{L}(x,q,\epsilon)$
@@ 17      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
18       /* Check whether $\epsilon$ satisfies controller’s privacy preference */
19       if $PRI_{min}/PRI_{max}\geq 1-(p/100)$ then
@@
Detailed Algorithm. Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ values is infinite, we set an upper bound and a range, e.g., 10 and 0.1, to ensure a finite set of options, e.g., 100 in this example. Then, for each $x$’s neighboring dataset $x_{-i}$, the algorithm computes $q(x_{-i})$ (line 1 - 1), which will be used to compute the PRI for each individual $i$. Next, the algorithm iterates over the set of $\epsilon$ in descending order (line 1). For each $\epsilon$, the algorithm computes the query output $r$ by applying the Laplace Mechanism $\mathcal{M}_{L}$ with parameter $\epsilon$ (line 1) and computes the corresponding PRI of each individual (line 1 - 1). Then, given the list of PRIs, the algorithm computes $PRI_{min}$ and $PRI_{max}$ (line 1). If $PRI_{min}/PRI_{max}\geq 1-(p/100)$, then the algorithm has found a suitable $\epsilon$ and a query output, so it stops and returns the output.
@@
Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRI profiles across executions of the mechanism; simultaneously, similar PRI profiles may be produced by different $\epsilon$ values. What is important is that the final query output the analyst releases has been subject to DP and the privacy loss of the individuals is consistent with the controller’s privacy preferences as expressed by $p$. In addition, we show empirically in Section 6.3.2 that stricter privacy preferences (i.e., smaller $p$) generally lead to smaller $\epsilon$. Figure 2 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs; for demonstration, we choose one of the five queries (Q5) we have run. As shown in the figure, as $p$ increases, the algorithm tends to choose larger $\epsilon$. Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
@@
Dataflows and leakage. It is critically important to understand what information is released to whom. Figure 3 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm. Individuals send their data to the controller; this dataflow remains unchanged from traditional central DP models (Near et al., 2021). The analyst submits a query to the controller. The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to find a suitable query output, which is an output produced by a standard DP mechanism (Laplace Mechanism in our work) with a value for $\epsilon$ selected by the algorithm. At this point, no information has been returned to the analyst yet. The PRI associated with each individual is never released to the analyst. The query output that satisfies the controller’s privacy preferences is only released with the controller’s approval. The value for $\epsilon$ that led to the query output is also not released by default.
Releasing $\epsilon$. Under certain scenarios (Dwork et al., 2019), $\epsilon$ may also need to be released. Since $\epsilon$ is chosen based on PRIs, which are data-dependent, the controller should release $\epsilon$ differential privately by incorporating techniques such as Exponential Mechanism (e.g., Report Noisy Max) (McSherry and Talwar, 2007) or Sparse Vector Technique (Dwork et al., 2009; Lyu et al., 2016) in the algorithm. Instead of choosing the largest $\epsilon$ that passes the PRI criterion, the algorithm will choose and return the first $\epsilon$ (which may or may not be the largest) that passes the PRI criterion while preserving DP.

### v3

@@ 4. Deriving Epsilon based on the PRI
In this section, we describe the intuition of using PRIs to express controllers’ privacy preferences (Section 4.1), then introduce an algorithm to choose $\epsilon$ by using PRIs as a more intuitive proxy (Section 4.2). We discuss the algorithm’s complexity and extensions in Section 4.3.
@@ 4.1. Using PRI to Express Privacy Preferences
Traditionally, controllers need to choose $\epsilon$ to convey their privacy preferences. However, choosing $\epsilon$ is challenging because it is hard to understand the specific effect of $\epsilon$ on the privacy of any one individual in the dataset. In contrast, the per-individual PRI indicates the realized privacy loss of an individual relative to other individuals in the dataset. The consequence is that the PRI becomes an actionable indicator that helps controllers decide whether a concrete query output matches their privacy preferences, in which case they may agree to release the output. Next, we convey the intuition on how the PRI is useful to controllers through a example.
Figure 2. Example of using PRI to find $\epsilon$. The query is to compute the average GPA of the three students. For each $\epsilon$ (including when no DP is used), we show a specific query output and the corresponding PRI of each student.
Example illustrating one use of the PRI. Figure 2 shows the resulting PRIs of three students after answering a query about their average GPA. If the query output were released without using DP (a), the analyst observing the output would learn much about those students’ grades, as indicated by the PRI (recall that the PRI is a proxy to privacy loss). If the controller used DP, they would have to choose $\epsilon$ and obtain the query output by applying a DP mechanism. This would lead to a noisy output and a PRI associated with each student (note that running the mechanism two times with the same $\epsilon$ does not guarantee the same PRIs are obtained). We observe that decreasing $\epsilon$ (b, c, d) results in noisier output, which causes the PRIs to become more similar across students. In the extreme, when the PRIs are equal (c), no student in the dataset is at a higher risk than others (they have the same privacy loss). Thus, between (b) and (c), the controller may be more comfortable releasing the specific output (average GPA of 3.275) produced by choosing $\epsilon=1$ because it protects each student equally, which may be a particular privacy preference the controller wants to achieve. Lastly, using a smaller $\epsilon=0.1$ (d) also makes the PRIs roughly similar, but at the cost of producing a query output with much worse accuracy than (c). One sensible choice is to choose the larger $\epsilon$ (c) as it yields higher utility for the analyst receiving the output.
The above is an illustrative example that does not work in practice, since finding an $\epsilon$ that makes the PRI exactly equal across individuals is generally infeasible: either such an $\epsilon$ does not exist, or it does, but it requires adding too much noise that the output offers little utility to the analyst.
@@ 4.2. The Find-$\epsilon$-from-PRI Algorithm
When a controller runs a query on a dataset with a DP mechanism, observing the PRIs associated with individuals in the dataset communicates to the controller the realized privacy loss of each individual compared to others if the query output is released. Controllers could then change $\epsilon$, rerun the DP mechanism, and observe the resulting PRIs: they could repeat this process until they find a profile of PRIs that represents their preferences. Rather than choosing $\epsilon$ and releasing the DP output, controllers first check the corresponding PRIs to understand how would releasing the output affect each individual’s privacy, and then they decide whether or not to release the output based on their privacy preferences on the individuals. We introduce an algorithm that implements this intuition, named Find-$\epsilon$-from-PRI. In Section 4.1, we have shown one criterion for choosing $\epsilon$ that leads to exactly equal PRIs, which is typically infeasible. Instead, we propose an alternative criterion to let controllers decide whether a query output respects their privacy preferences. We explain the key idea and then the algorithm.
Key Idea. We let controllers express their privacy preferences by indicating a cutoff for the highest privacy loss of any individual in the dataset. This is done by requesting exactly one parameter from the controllers, a percentage $p$ ($0\leq p\leq 100$), which is easier to interpret than $\epsilon$. We compute the minimum and maximum PRI, $PRI_{min}$ and $PRI_{max}$. Then we find an $\epsilon$ that produces a DP output with its corresponding PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$. In other words, we let controllers specify that the maximum PRI should deviate no more than $p$ percent from the minimum PRI. For instance, setting $p=0$ means the PRIs should be exactly equal.
Unlike $\epsilon$, the parameter $p$ directly conveys how close the individuals’ realized privacy loss should be — the PRI is a relative indicator of ex-post privacy loss for each individual. Smaller $p$ means more privacy at the cost of worse accuracy, hence controllers can trade privacy for accuracy by setting a larger $p$. The algorithm aims to find the largest $\epsilon$ that produces a list of PRIs such that $PRI_{min}/PRI_{max}\geq 1-(p/100)$, because if controllers’ privacy preferences can be satisfied, larger $\epsilon$ is preferred as it leads to more accurate output.
Input : DP mechanism $\mathcal{M}$; dataset $x=(x_{1},\cdots,x_{n})$; query $q$; percentage $p$
Output : Query output $r$
1
2/* Initialize set of $\epsilon$ values to test */
@@ 3 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
4
5/* Compute $q(x_{-i})$ for all individual $i$ using memoization */
@@ 9             $C[x_{i}]\leftarrow q(x_{-i})$
15       $r\leftarrow\mathcal{M}(x,q,\epsilon,\delta=\frac{1}{n^{2}})$
16
@@ 21                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
22            if $\mathcal{M}$ is ’Gaussian’ then
23                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{2}$
@@ 26      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
27
28      /* Check whether PRIs satisfy controller’s privacy preference */
29       if $PRI_{min}/PRI_{max}\geq 1-\frac{p}{100}$ then
@@
Detailed Algorithm. Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output $r$. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ values is infinite, we set an upper bound and a range, e.g., 10 and 0.1, to ensure a finite set of options, e.g., 100 in this example. Then, for each $x$’s neighboring dataset $x_{-i}$, the algorithm computes $q(x_{-i})$ (line 1 - 1), which will be used to compute the PRI for each individual $i$. Next, the algorithm iterates over the set of $\epsilon$ in descending order (line 1). For each $\epsilon$, the algorithm computes the query output $r$ by applying the given DP Mechanism $\mathcal{M}$ with parameter $\epsilon$ and a fixed $\delta$ (line 1), then computes the corresponding PRI of each individual (line 1 - 1). Then, given the list of PRIs, the algorithm computes $PRI_{min}$ and $PRI_{max}$ (line 1). If $PRI_{min}/PRI_{max}\geq 1-(p/100)$, then the algorithm has found a suitable $\epsilon$ and a query output, so it stops and returns the output. Note that at this point $\epsilon$ is not released to the analysts, and only the controller knows which $\epsilon$ is chosen by the algorithm.
@@
Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRIs across executions of the mechanism; simultaneously, similar PRIs may be produced by different $\epsilon$ values. What is important is that the final query output the analyst releases has been subject to DP and the privacy loss of the individuals is consistent with the controller’s privacy preferences as expressed by $p$.
In addition, we show empirically in Section 6.3.1 that stricter privacy preferences (i.e., smaller $p$) lead to smaller $\epsilon$, as expected. For demonstration, we present one of our evaluation results (Q4 in Figure 10). Figure 3 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs. As shown in the figure, as $p$ increases, the algorithm tends to choose larger $\epsilon$. Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 0.1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
@@
Dataflows and leakage. It is critically important to understand what information is released to whom. Figure 4 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm to find $\epsilon$ and release query output to the analyst. Individuals send their data to the controller; this dataflow remains unchanged from the traditional central DP model (Near et al., 2021). The analyst submits a query to the controller. The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to release a suitable query output, which is an output produced by a DP mechanism (Laplace Mechanism or Gaussian Mechanism in our work) with an $\epsilon$ selected by the algorithm. At this point, no information has been returned to the analyst yet. The PRI associated with each individual is never released to the analyst. The query output that satisfies the controller’s privacy preferences is only released with the controller’s approval. The value for $\epsilon$ that led to the query output is also not released since $\epsilon$ is chosen based on PRIs, which are data-dependent, so releasing it directly might jeopardize the DP guarantee. In Section 5, we explain a modification of the algorithm that permits releasing $\epsilon$ differential privately.
Claim: Algorithm 1 is $(\epsilon,\delta)$-DP, where $\epsilon$ is the value found by the algorithm that satisfies the PRI threshold, or 0 if no suitable $\epsilon$ is found.
Algorithm 1 only releases the query output $r$, which is computed using an $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when $\mathcal{M}$ is Laplace Mechanism and $\delta=\frac{1}{n^{2}}$ when $\mathcal{M}$ is Gaussian Mechanism). If no suitable $\epsilon$ is found, the algorithm does not release $r$. Therefore, the algorithm satisfies $(\epsilon,\delta)$-DP.
4.3. Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel. Thus, we implement parallelism and compute the query output for a chunk of records in each process, and we show in Section 6.2 that the algorithm is both efficient and scalable.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP. The PRI complements these existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
5. Releasing $\epsilon$ and Bounding Privacy Loss
In Section 5.1, we propose a modification of Find-$\epsilon$-from-PRI that finds $\epsilon$ and releases both the query output and $\epsilon$ simultaneously while preserving DP. In Section 5.2, we discuss how to bound the total privacy loss in the presence of multiple queries and we provide the end-to-end DP guarantee in Section 5.3.
5.1. Privately Releasing $\epsilon$
Since Algorithm 1 finds $\epsilon$ in a data-dependent manner, releasing the chosen $\epsilon$ directly may reveal additional information about the dataset and violate the DP guarantees. Therefore, a modification of the algorithm is needed to ensure DP is preserved if the controller wants to release both the query output $r$ and $\epsilon$ to the analysts.
Key Idea. Our key idea is to leverage the Sparse Vector Technique (SVT) (Dwork et al., 2009) to test multiple $\epsilon$ values (from largest to smallest) and release the first $\epsilon$ that satisfies the PRI threshold specified by the controller. Algorithm 1 tests the set of $\epsilon$ values and chooses the largest $\epsilon$ that passes the PRI threshold; by incorporating SVT, the modified algorithm chooses approximately the largest $\epsilon$ that passes the threshold so releasing it preserves DP.
Input : Dataset $x$; a stream of SVT queries $\mathcal{Q}_{svt}=q_{1},q_{2}\cdots$ with global sensitivity at most $\Delta_{svt}$; threshold $\tau$; privacy budget $\epsilon_{svt}^{\prime}$ and $\epsilon_{svt}^{\prime\prime}$
Output : A stream of answers $a_{1},a_{2}\cdots$ where each $a_{i}\in\{\top,\bot\}$
1
[…]
3for $q_{i}\in\mathcal{Q}_{svt}$ do
4
5      if $q_{i}(x)+Lap\left(\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}\right)\geq\tau+\rho$ then
6             Output $a_{i}=\top$
7             return
8      else
9             Output $a_{i}=\bot$
10
Algorithm 2 SVT algorithm proposed by Lyu et al. (Lyu et al., 2016)
The SVT algorithm. Algorithm 2 shows the standard SVT algorithm proposed by  Lyu et al. (Lyu et al., 2016)  [^4] The algorithm from  (Lyu et al., 2016) has an additional parameter $c$ (the maximum number of queries the algorithm can answer $\top$). We set $c=1$, since we only need one positive answer for our modified algorithm., which we will adopt and incorporate into our modified algorithm to release $\epsilon$ privately. SVT takes a dataset, a stream of queries (each with a global sensitivity of at most $\Delta_{svt}$), a given threshold, and a privacy budget $\epsilon_{svt}$ that is divided into $\epsilon_{svt}^{\prime}$ and $\epsilon_{svt}^{\prime\prime}$ (i.e., $\epsilon_{svt}=\epsilon_{svt}^{\prime}+\epsilon_{svt}^{\prime\prime}$), then outputs a stream of answers indicating whether each query output is above or below the threshold. The algorithm works by first initializing a Laplacian noise ($\rho$) scaled to $\frac{\Delta_{svt}}{\epsilon_{svt}^{\prime}}$ (line 2). For each SVT query, it generates another term of Laplacian noise scaled to $\frac{2\Delta_{svt}}{\epsilon_{svt}^{\prime\prime}}$ and adds the noise to the query output. Then the algorithm compares the noisy query output with the noisy threshold (line 2); if it is above or equal to the threshold, the algorithm outputs $\top$ indicating the query output approximately passes the threshold, otherwise, it outputs $\bot$. Note that the algorithm answers the SVT queries in a streaming fashion. Overall, the algorithm satisfies $\epsilon_{svt}$-DP.
Input : DP mechanism $\mathcal{M}$; dataset $x=(x_{1},\cdots,x_{n})$; query $q$; variance threshold $\tau$; SVT budget $\epsilon_{svt}$
Output : Query output $r$, chosen $\epsilon$
1
2/* Assign privacy budget for running SVT */
3 $\epsilon_{svt}^{\prime},\epsilon_{svt}^{\prime\prime}\leftarrow\frac{\epsilon_{svt}}{1+2^{2/3}},\frac{2^{2/3}\epsilon_{svt}}{1+2^{2/3}}$
4
5/* Global sensitivity of variance */
6 $\Delta\leftarrow\frac{1}{n}$
7
8/* Generate noise to add to PRI threshold */
9 $\rho=Lap\left(\frac{\Delta}{\epsilon_{svt}^{\prime}}\right)$
10
11/* Initialize set of $\epsilon$ values to test */
12 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
13
14/* Compute $q(x_{-i})$ for each individual $i$ using memoization */
15 $C\leftarrow map()$
16 for $x_{i}\in x$ do
17       if $x_{i}\notin C$ then
18             $C[x_{i}]\leftarrow q(x_{-i})$
19
20
21for $\epsilon\in\mathcal{E}$ do
22
23      /* Compute query output using DP with current $\epsilon$/
24       $r\leftarrow\mathcal{M}(x,q,\epsilon,\delta=\frac{1}{n^{2}})$
25
26      /* Compute corresponding PRI of each individual */
27       $PRIs\leftarrow list()$
28       for $x_{i}\in C$ do
29             if $\mathcal{M}$ is ’Laplace’ then
30                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
31            if $\mathcal{M}$ is ’Gaussian’ then
32                  $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{2}$
33            $PRIs.add(PRI_{i})$
34
35      /* Normalize PRIs and compute variance */
36       $PRI_{max}\leftarrow\max(PRIs)$
37       for $i\in range(PRIs.length)$ do
38             $PRIs[i]\leftarrow\frac{PRIs[i]}{PRI_{max}}$
39      $Var\leftarrow$ compute-variance($PRIs$)
40
41      /* Using SVT to release the output $r$ and $\epsilon$ that approximately passes variance threshold */
42
43      if $-Var+Lap\left(\frac{2\Delta}{\epsilon_{svt}^{\prime\prime}}\right)\geq-\tau+\rho$ then
44             return $r,\epsilon$
45
46
return nil
Algorithm 3 Find-and-release-$\epsilon$-from-PRI
What is a good SVT query for choosing $\epsilon$? To incorporate SVT in our algorithm, each SVT query corresponds to a specific $\epsilon$ the algorithm is testing, and each query should output a value based on PRIs computed from the $\epsilon$; naturally, this value could be the existing PRI criterion we already have in Algorithm 1, which is $\frac{PRI_{min}}{PRI_{max}}$. Then the algorithm applies SVT by comparing the PRI criterion to the threshold (i.e., the percentage $p$ specified by controllers) and releasing $\epsilon$ of which the corresponding PRI criterion approximately passes the threshold. However, the current PRI criterion has one caveat that makes it difficult to incorporate with SVT: the SVT query, which is a function that outputs $\frac{PRI_{min}}{PRI_{max}}$, has a sensitivity $\Delta_{svt}=1$. This means the algorithm will be adding a Laplacian noise scaled to $\frac{2}{\epsilon_{svt}^{\prime\prime}}$ to $\frac{PRI_{min}}{PRI_{max}}$; clearly, when $\epsilon_{svt}^{\prime\prime}$ is small, the magnitude of added noise will be too large and make the PRI criterion useless since it will be too noisy to choose a reasonable $\epsilon$.
Using variance as the new PRI criterion. To address this caveat, we opt for a different PRI criterion when using SVT — by using the variance of normalized PRIs. Each SVT query will output the variance of PRIs for the corresponding $\epsilon$, and the variance is compared to a threshold specified by the controller; if the variance is approximately smaller than the threshold, then a suitable $\epsilon$ is found. The normalized PRIs will have values between 0 and 1 — this still preserves the PRI’s meaning as a relative indicator of ex-post per-individual privacy loss. The variance of the PRIs will have a sensitivity of $\frac{1}{n}$, where n is the number of records in the dataset; this means when $n$ is relatively large, the sensitivity of the variance (hence the Laplacian noise added to the variance) will be small! In addition, the variance of PRIs shares the same intuitiveness as $\frac{PRI_{min}}{PRI_{max}}$: if the controllers want the individuals to have similar realized privacy loss (hence more protected), then they should set a lower variance threshold, and vice versa.
The Find-and-release-$\epsilon$-from-PRI algorithm. We present our modified Find-and-release-$\epsilon$-from-PRI algorithm in Algorithm 3. The main modification lies in line 3 -  3 and line 3 -  3. In line 3, the algorithm allocate the SVT budget $\epsilon_{svt}$ to have $\epsilon_{svt}^{\prime}:\epsilon_{svt}^{\prime\prime}=1:2^{\frac{2}{3}}$; we choose this allocation strategy from  (Lyu et al., 2016) as it leads to better accuracy of SVT. Then the algorithm iterates over each $\epsilon$ and computes its corresponding PRIs the same way as in Algorithm 1 (line 3 -  3). It then normalizes each PRI and computes the variance (line 3), and lastly, it compares the noisy variance with the noisy threshold (controllers want $Var\leq\tau$, which is equivalent to $-Var\geq-\tau$, and returns the current output $r$ and $\epsilon$ if it passes the threshold.
The main advantage of using SVT is that it always consumes a fixed budget $\epsilon_{svt}$ regardless of the number of SVT queries. Controllers understand the effect of using a specific $\epsilon_{svt}$ and running SVT to release $\epsilon$, since they know exactly how much noise is added to both the variance and the variance threshold $\tau$ they have specified. They can decide whether they are comfortable releasing an output that does not exactly conform to their privacy preferences; if not, they may use a different $\epsilon_{svt}$ or $\tau$ and rerun the algorithm.
Claim: Algorithm 3 is $(\epsilon+\epsilon_{svt},\delta)$-DP, where $\epsilon$ is the value released using SVT, or 0 if no suitable $\epsilon$ is found.
Algorithm 3 can be divided into two steps. In the first step, it uses a $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when the mechanism is Laplace) to compute the query output $r$. In the second step, it uses SVT, which is $\epsilon_{svt}$-DP, to output whether the SVT query for the given $\epsilon,r$ is above the threshold, and releases $\epsilon,r$ if it is above the threshold. Therefore, by sequential composition (Dwork et al., 2014), Algorithm 3 satisfies $(\epsilon+\epsilon_{svt},\delta)$-DP.